# On Random Fields Associated with Analytic Wavelet Transform

### v1

@@ 5.12. Proof of Corollary 5
[…]
$dr_{1}dr_{2}d\theta_{1}d\theta_{2}$
@@
$\times\textup{exp}\left(-\frac{\Gamma_{11}r_{1}^{2}+\Gamma_{22}r_{2}^{2}-2\Re(e^{i(\theta_{1}-\theta_{2})}\overline{\Gamma_{12}})r_{1}r_{2}}{\textup{det}(\Gamma)}\right)$
$dr_{1}dr_{2}d\theta_{1}d\theta_{2}$
@@
$\textup{exp}\left(-r_{1}^{2}-r_{2}^{2}+\frac{2|\Gamma_{12}|}{\sqrt{\Gamma_{11}\Gamma_{22}}}\cos\Big{(}\theta_{1}-\theta_{2}-\arg(\Gamma_{12})\Big{)}r_{1}r_{2}\right)$
(109) $dr_{1}dr_{2}d\theta_{1}d\theta_{2}.$
@@
(110) $=$ $h(\frac{|\Gamma_{12}|}{\sqrt{\Gamma_{11}\Gamma_{22}}},\arg(\Gamma_{12})),$
@@
$\textup{exp}\left(-r_{1}^{2}-r_{2}^{2}\right)\textup{exp}\Big{(}2u\cos\big{(}\theta_{1}-\theta_{2}-v\big{)}r_{1}r_{2}\Big{)}$
$dr_{1}dr_{2}d\theta_{1}d\theta_{2}$
@@
$\underset{u\rightarrow 0+}{\lim}\frac{h(u,v)}{u}$
$=$ $\underset{u\rightarrow 0+}{\lim}-2u\widetilde{h}(u,v)$
@@
$=$ $\frac{2}{\pi^{2}}\left(\int_{0}^{\infty}r^{2}e^{-r^{2}}dr\right)^{2}$
$\times\int_{0}^{2\pi}\int_{0}^{2\pi}\left(\theta_{1}-\pi\right)\left(\theta_{2}-\pi\right)\cos\big{(}\theta_{1}-\theta_{2}-v\big{)}d\theta_{1}d\theta_{2}$
@@ 5.13. Proof of Corollary 6
$=$ $\int_{0}^{\infty}\int_{0}^{2\pi}e^{i\theta}p\left(re^{i\theta};W_{\Phi}(t_{1},s_{1})\overline{W_{\Phi}(t_{2},s_{2})}\right)rd\theta dr,$
@@
(115) $=$ $\frac{1}{\pi^{2}\textup{det}(\Gamma)|w_{1}|^{2}}\textup{exp}\left(-\left[\begin{array}[]{cc}w_{1}\\ \overline{z}/\overline{w_{1}}\end{array}\right]^{*}\right$.
(118) $\left.\Gamma^{-1}\left[\begin{array}[]{cc}w_{1}\\ \overline{z}/\overline{w_{1}}\end{array}\right]\right),\ w_{1},q\in\mathbb{C},$
@@
(134) $\left[\begin{array}[]{cc}r\\ \overline{z}r^{-1}\end{array}\right]^{*}\Gamma^{-1}\left[\begin{array}[]{cc}r\\ \overline{z}r^{-1}\end{array}\right]$
(135) $=$ $\frac{\Gamma_{22}r^{2}+\Gamma_{11}|z|^{2}r^{-2}-2\mathfrak{R}\left(\overline{z}\Gamma_{12}\right)}{\det(\Gamma)}.$
@@
$=$ $\textup{exp}\left(\frac{2\mathfrak{R}\left(\overline{z}\Gamma_{12}\right)}{\det(\Gamma)}\right)$
$\times\int_{0}^{\infty}\frac{2}{\pi\textup{det}(\Gamma)r}\textup{exp}\left(-\frac{\Gamma_{22}r^{2}+\Gamma_{11}|z|^{2}r^{-2}}{\det(\Gamma)}\right)dr.$
@@
$\int_{0}^{\infty}\frac{2}{\pi\textup{det}(\Gamma)r}\textup{exp}\left(-\frac{\Gamma_{22}r^{2}+\Gamma_{11}|z|^{2}r^{-2}}{\det(\Gamma)}\right)dr$
(137) $=$ $\frac{2}{\pi\textup{det}(\Gamma)}K_{0}\left(\frac{2\sqrt{\Gamma_{11}\Gamma_{22}}}{\det(\Gamma)}|z|\right).$
@@
(138) $p\left(z;W_{\Phi}(t_{1},s_{1})\overline{W_{\Phi}(t_{2},s_{2})}\right)$
$=$ $\frac{2}{\pi\textup{det}(\Gamma)}\textup{exp}\left(\frac{2\mathfrak{R}\left(\overline{z}\Gamma_{12}\right)}{\det(\Gamma)}\right)K_{0}\left(\frac{2\sqrt{\Gamma_{11}\Gamma_{22}}}{\det(\Gamma)}|z|\right)$
@@
$=$ $\int_{0}^{\infty}\int_{0}^{2\pi}e^{i\theta}\left[\frac{2}{\pi\textup{det}(\Gamma)}\textup{exp}\left(\frac{2\mathfrak{R}\left(re^{-i\theta}\Gamma_{12}\right)}{\det(\Gamma)}\right)\right$.
(139) $\left.K_{0}\left(\frac{2\sqrt{\Gamma_{11}\Gamma_{22}}}{\det(\Gamma)}r\right)\right]rd\theta dr.$
@@
$=$ $\frac{2}{\pi\textup{det}(\Gamma)}\left(2\pi e^{i\arg\left(\Gamma_{12}\right)}\right)\int_{0}^{\infty}I_{1}\left(\frac{2|\Gamma_{12}|}{\det(\Gamma)}r\right)$
$K_{0}\left(\frac{2\sqrt{\Gamma_{11}\Gamma_{22}}}{\det(\Gamma)}r\right)rdr$
$=$ $\frac{4e^{i\arg\left(\Gamma_{12}\right)}}{\textup{det}(\Gamma)}\left(\frac{\det(\Gamma)}{2\sqrt{\Gamma_{11}\Gamma_{22}}}\right)^{2}\int_{0}^{\infty}I_{1}\left(\frac{|\Gamma_{12}|}{\sqrt{\Gamma_{11}\Gamma_{22}}}u\right)$
$K_{0}\left(u\right)udu$
@@
$\int_{0}^{\infty}I_{1}\left(cu\right)K_{0}\left(u\right)udu$
$=$ $\int_{0}^{\infty}\left[\frac{1}{2}cu\overset{\infty}{\underset{k=0}{\sum}}\frac{\left(\frac{1}{4}c^{2}u^{2}\right)^{k}}{k!\Gamma(k+2)}\right]K_{0}\left(u\right)udu$
@@
$=$ $\frac{\pi}{4}e^{i\arg\left(\Gamma_{12}\right)}\frac{\det(\Gamma)}{\Gamma_{11}\Gamma_{22}}\frac{|\Gamma_{12}|}{\sqrt{\Gamma_{11}\Gamma_{22}}}{}_{2}F_{1}\left(\frac{3}{2},\frac{3}{2};2;\frac{|\Gamma_{12}|^{2}}{\Gamma_{11}\Gamma_{22}}\right)$
@@ 5.14. Proof of Corollary 7
$=$ $\int_{0}^{\infty}\int_{0}^{2\pi}p\left(r_{1},r_{2},\theta_{1},\theta_{2};|W_{\Phi}(t_{1},s_{1})|,|W_{\Phi}(t_{2},s_{2})|,\right$.
$\left.\Theta_{\Phi}(t_{1},s_{1}),\Theta_{\Phi}(t_{2},s_{2})\right)d\phi d\rho$
@@
$=$ $\int_{0}^{\infty}\int_{0}^{2\pi}\frac{r_{1}r_{2}}{\pi^{2}\textup{det}(\Gamma)}\textup{exp}\left(-\frac{\Gamma_{22}r_{1}^{2}+\Gamma_{11}r_{2}^{2}}{\textup{det}(\Gamma)}\right)$
(145) $\times\textup{exp}\left(\frac{2r_{1}r_{2}\Re(e^{i(\theta_{1}-\theta_{2})}\overline{\Gamma_{12}})}{\textup{det}(\Gamma)}\right)d\theta_{1}dr_{2}.$
@@
$=$ $p\left(r_{1};|W_{\Phi}(t_{1},s_{1})|\right)p\left(\theta_{2};\Theta_{\Phi}(t_{2},s_{2})\right),$

### v2

@@ 5.12. Proof of Corollary 5
$\times\textup{exp}\left(-\begin{bmatrix}r_{1}e^{i\theta_{1}}\\ r_{2}e^{i\theta_{2}}\end{bmatrix}^{*}\Gamma^{-1}\begin{bmatrix}r_{1}e^{i\theta_{1}}\\ r_{2}e^{i\theta_{2}}\end{bmatrix}\right)dr_{1}dr_{2}d\theta_{1}d\theta_{2}$
@@
$\times\textup{exp}\left(-\frac{\Gamma_{11}r_{1}^{2}+\Gamma_{22}r_{2}^{2}-2\Re(e^{i(\theta_{1}-\theta_{2})}\overline{\Gamma_{12}})r_{1}r_{2}}{\textup{det}(\Gamma)}\right)dr_{1}dr_{2}d\theta_{1}d\theta_{2}$
@@
(111) $\textup{exp}\left(-r_{1}^{2}-r_{2}^{2}+\frac{2|\Gamma_{12}|}{\sqrt{\Gamma_{11}\Gamma_{22}}}\cos\Big(\theta_{1}-\theta_{2}-\arg(\Gamma_{12})\Big)r_{1}r_{2}\right)dr_{1}dr_{2}d\theta_{1}d\theta_{2}.$
@@
(112) $\textup{Cov}\left(\Theta_{\Phi}(t_{1},s_{1}),\Theta_{\Phi}(t_{2},s_{2})\right)=h(\frac{|\Gamma_{12}|}{\sqrt{\Gamma_{11}\Gamma_{22}}},\arg(\Gamma_{12})),$
@@
$\textup{exp}\left(-r_{1}^{2}-r_{2}^{2}\right)\textup{exp}\Big(2u\cos\big(\theta_{1}-\theta_{2}-v\big)r_{1}r_{2}\Big)dr_{1}dr_{2}d\theta_{1}d\theta_{2}$
@@
$\underset{u\rightarrow 0+}{\lim}\frac{h(u,v)}{u}=$ $\underset{u\rightarrow 0+}{\lim}-2u\widetilde{h}(u,v)$
@@
$=$ $\frac{2}{\pi^{2}}\left(\int_{0}^{\infty}r^{2}e^{-r^{2}}dr\right)^{2}\int_{0}^{2\pi}\int_{0}^{2\pi}\left(\theta_{1}-\pi\right)\left(\theta_{2}-\pi\right)\cos\big(\theta_{1}-\theta_{2}-v\big)d\theta_{1}d\theta_{2}$
@@ 5.13. Proof of Corollary 6
(114) $\mathbb{E}\left[e^{i(\Theta_{\Phi}(t_{1},s_{1})-\Theta_{\Phi}(t_{2},s_{2}))}\right]=\int_{0}^{\infty}\int_{0}^{2\pi}\hskip-7.11317pte^{i\theta}p\left(re^{i\theta};W_{\Phi}(t_{1},s_{1})\overline{W_{\Phi}(t_{2},s_{2})}\right)rd\theta dr,$
@@
(119) $=$ $\frac{1}{\pi^{2}\textup{det}(\Gamma)|w_{1}|^{2}}\textup{exp}\left(-\left[\begin{array}[]{cc}w_{1}\\ \overline{z}/\overline{w_{1}}\end{array}\right]^{*}\Gamma^{-1}\left[\begin{array}[]{cc}w_{1}\\ \overline{z}/\overline{w_{1}}\end{array}\right]\right),\ w_{1},q\in\mathbb{C},$
@@
(135) $\left[\begin{array}[]{cc}r\\ \overline{z}r^{-1}\end{array}\right]^{*}\Gamma^{-1}\left[\begin{array}[]{cc}r\\ \overline{z}r^{-1}\end{array}\right]=\frac{\Gamma_{22}r^{2}+\Gamma_{11}|z|^{2}r^{-2}-2\mathfrak{R}\left(\overline{z}\Gamma_{12}\right)}{\det(\Gamma)}.$
@@
$=$ $\textup{exp}\left(\frac{2\mathfrak{R}\left(\overline{z}\Gamma_{12}\right)}{\det(\Gamma)}\right)\int_{0}^{\infty}\frac{2}{\pi\textup{det}(\Gamma)r}\textup{exp}\left(-\frac{\Gamma_{22}r^{2}+\Gamma_{11}|z|^{2}r^{-2}}{\det(\Gamma)}\right)dr.$
@@
(137) $\int_{0}^{\infty}\frac{2}{\pi\textup{det}(\Gamma)r}\textup{exp}\left(-\frac{\Gamma_{22}r^{2}+\Gamma_{11}|z|^{2}r^{-2}}{\det(\Gamma)}\right)dr=\frac{2}{\pi\textup{det}(\Gamma)}K_{0}\left(\frac{2\sqrt{\Gamma_{11}\Gamma_{22}}}{\det(\Gamma)}|z|\right).$
@@
(138) $p\left(z;W_{\Phi}(t_{1},s_{1})\overline{W_{\Phi}(t_{2},s_{2})}\right)=\frac{2}{\pi\textup{det}(\Gamma)}\textup{exp}\left(\frac{2\mathfrak{R}\left(\overline{z}\Gamma_{12}\right)}{\det(\Gamma)}\right)K_{0}\left(\frac{2\sqrt{\Gamma_{11}\Gamma_{22}}}{\det(\Gamma)}|z|\right)$
@@
(139) $=$ $\int_{0}^{\infty}\int_{0}^{2\pi}e^{i\theta}\left[\frac{2}{\pi\textup{det}(\Gamma)}\textup{exp}\left(\frac{2\mathfrak{R}\left(re^{-i\theta}\Gamma_{12}\right)}{\det(\Gamma)}\right)K_{0}\left(\frac{2\sqrt{\Gamma_{11}\Gamma_{22}}}{\det(\Gamma)}r\right)\right]rd\theta dr.$
@@
$=$ $\frac{2}{\pi\textup{det}(\Gamma)}\left(2\pi e^{i\arg\left(\Gamma_{12}\right)}\right)\int_{0}^{\infty}I_{1}\left(\frac{2|\Gamma_{12}|}{\det(\Gamma)}r\right)K_{0}\left(\frac{2\sqrt{\Gamma_{11}\Gamma_{22}}}{\det(\Gamma)}r\right)rdr$
$=$ $\frac{4e^{i\arg\left(\Gamma_{12}\right)}}{\textup{det}(\Gamma)}\left(\frac{\det(\Gamma)}{2\sqrt{\Gamma_{11}\Gamma_{22}}}\right)^{2}\int_{0}^{\infty}I_{1}\left(\frac{|\Gamma_{12}|}{\sqrt{\Gamma_{11}\Gamma_{22}}}u\right)K_{0}\left(u\right)udu$
@@
$\int_{0}^{\infty}I_{1}\left(cu\right)K_{0}\left(u\right)udu=$ $\int_{0}^{\infty}\left[\frac{1}{2}cu\overset{\infty}{\underset{k=0}{\sum}}\frac{\left(\frac{1}{4}c^{2}u^{2}\right)^{k}}{k!\Gamma(k+2)}\right]K_{0}\left(u\right)udu$
@@
$\mathbb{E}\left[e^{i(\Theta_{\Phi}(t_{1},s_{1})-\Theta_{\Phi}(t_{2},s_{2}))}\right]=$ $\frac{\pi}{4}e^{i\arg\left(\Gamma_{12}\right)}\frac{\det(\Gamma)}{\Gamma_{11}\Gamma_{22}}\frac{|\Gamma_{12}|}{\sqrt{\Gamma_{11}\Gamma_{22}}}{}_{2}F_{1}\left(\frac{3}{2},\frac{3}{2};2;\frac{|\Gamma_{12}|^{2}}{\Gamma_{11}\Gamma_{22}}\right)$
@@ 5.14. Proof of Corollary 7
(142) $p\left(r_{1},\theta_{2};|W_{\Phi}(t_{1},s_{1})|,\Theta_{\Phi}(t_{2},s_{2})\right)$
$=$ $\int_{0}^{\infty}\int_{0}^{2\pi}p\left(r_{1},r_{2},\theta_{1},\theta_{2};|W_{\Phi}(t_{1},s_{1})|,|W_{\Phi}(t_{2},s_{2})|,\Theta_{\Phi}(t_{1},s_{1}),\Theta_{\Phi}(t_{2},s_{2})\right)d\phi d\rho$
@@
$=$ $\int_{0}^{\infty}\int_{0}^{2\pi}\frac{r_{1}r_{2}}{\pi^{2}\textup{det}(\Gamma)}\textup{exp}\left(-\frac{\Gamma_{22}r_{1}^{2}+\Gamma_{11}r_{2}^{2}}{\textup{det}(\Gamma)}\right)\textup{exp}\left(\frac{2r_{1}r_{2}\Re(e^{i(\theta_{1}-\theta_{2})}\overline{\Gamma_{12}})}{\textup{det}(\Gamma)}\right)d\theta_{1}dr_{2}.$
@@
$p\left(r_{1},\theta_{2};|W_{\Phi}(t_{1},s_{1})|,\Theta_{\Phi}(t_{2},s_{2})\right)=p\left(r_{1};|W_{\Phi}(t_{1},s_{1})|\right)p\left(\theta_{2};\Theta_{\Phi}(t_{2},s_{2})\right),$
@@
5.15. Proof of Proposition 8
By the Skorohod representation theorem [40] (see also [41, Theorem 3.1] and [33, Theorem 3.1]), on a sufficiently rich probability space, there exist random variables $\{\widehat{\epsilon}_{j}\}_{j\in\mathbb{Z}}$ and mean-zero independent Gaussian random variables $\{z_{j}\}_{j\in\mathbb{Z}}$, defined on the same probability space, such that $\{\widehat{\epsilon}_{j}\}_{j\in\mathbb{Z}}$ has the same distribution as $\{\epsilon_{j}\}_{j\in\mathbb{Z}}$ and $\textup{Var}(z_{j})=\textup{Var}(\epsilon_{j})$ for all $j\in\mathbb{Z}$. Let $\widetilde{W}_{\widehat{\epsilon}}$ denote the discretized AWT of $\{\widehat{\epsilon}_{j}\}_{j\in\mathbb{Z}}$, and denote the probability measure of $\{\widetilde{W}_{\widehat{\epsilon}}(t_{i},s_{\ell})\mid 1\leq i\leq n,\ 1\leq\ell\leq d\}$ by $\nu_{\widehat{\epsilon}}^{(n)}$. We have
(146) $\nu_{\epsilon}^{(n)}=\nu_{\widehat{\epsilon}}^{(n)}.$
Define
$I_{n}=[-\lceil\sqrt{n}As_{d}\rceil+1,\lceil\sqrt{n}As_{d}\rceil+n]\cap\mathbb{Z}.$
For $h\in I_{n}$, define the partial sum
[…]
with $S_{\widehat{\epsilon}}(h)=0$ when $h\leq-\lceil\sqrt{n}As_{d}\rceil$. By (48) and the summation by parts, for any $1\leq i\leq n$ and $1\leq\ell\leq d$,
$\widetilde{W}_{\widehat{\epsilon}}(t_{i},s_{\ell})=$ $\frac{1}{n^{1/4}}\underset{j\in D^{(\ell)}}{\sum}\left(S_{\widehat{\epsilon}}(i+j)-S_{\widehat{\epsilon}}(i+j-1)\right)\overline{\psi_{j}^{(\ell)}}$
$=$ $\frac{1}{n^{1/4}}S_{\widehat{\epsilon}}\left(i+\lceil\sqrt{n}As_{\ell}\rceil\right)\overline{\psi_{\lceil\sqrt{n}As_{\ell}\rceil+1}^{(\ell)}}$
$-\frac{1}{n^{1/4}}S_{\widehat{\epsilon}}\left(i-\lceil\sqrt{n}As_{\ell}\rceil-1\right)\overline{\psi_{-\lceil\sqrt{n}As_{\ell}\rceil}^{(\ell)}}$
[…]
Similarly, the discretized AWT of the Gaussian sequence $\{z_{j}\}_{j\in\mathbb{Z}}$ can be expressed as
$\hskip-8.5359pt\widetilde{W}_{z}(t_{i},s_{\ell})=$ $\frac{1}{n^{1/4}}S_{z}\left(i+\lceil\sqrt{n}As_{\ell}\rceil\right)\overline{\psi_{\lceil\sqrt{n}As_{\ell}\rceil+1}^{(\ell)}}$
$-\frac{1}{n^{1/4}}S_{z}\left(i-\lceil\sqrt{n}As_{\ell}\rceil-1\right)\overline{\psi_{-\lceil\sqrt{n}As_{\ell}\rceil}^{(\ell)}}$
[…]
where
[…]
and $S_{z}(h)=0$ for $h\leq-\lceil\sqrt{n}As_{d}\rceil$. By (147) and (148),
$n^{1/4}\left(\widetilde{W}_{\widehat{\epsilon}}(t_{i},s_{\ell})-\widetilde{W}_{z}(t_{i},s_{\ell})\right)$
$=$ $\left[S_{\widehat{\epsilon}}\left(i+\lceil\sqrt{n}As_{\ell}\rceil\right)-S_{z}\left(i+\lceil\sqrt{n}As_{\ell}\rceil\right)\right]\overline{\psi_{\lceil\sqrt{n}As_{\ell}\rceil+1}^{(\ell)}}$
$-\left[S_{\widehat{\epsilon}}\left(i-\hskip-2.84544pt\lceil\sqrt{n}As_{\ell}\rceil\hskip-2.84544pt-\hskip-1.42271pt1\right)-S_{z}\left(i-\hskip-2.84544pt\lceil\sqrt{n}As_{\ell}\rceil\hskip-2.84544pt-\hskip-1.42271pt1\right)\right]\hskip-1.42271pt\overline{\psi_{-\lceil\sqrt{n}As_{\ell}\rceil}^{(\ell)}}$
[…]
From (49), both $|\psi_{\lceil\sqrt{n}As_{\ell}\rceil+1}^{(\ell)}|$ and $|\psi_{-\lceil\sqrt{n}As_{\ell}\rceil}^{(\ell)}|$ are bounded by $n^{-1/4}s_{\ell}^{-1/2}\|\psi\|_{L^{\infty}}$. Hence,
$\left|\widetilde{W}_{\widehat{\epsilon}}(t_{i},s_{\ell})-\widetilde{W}_{z}(t_{i},s_{\ell})\right|\leq$ $\frac{\|\psi\|_{L^{\infty}}}{\sqrt{ns_{\ell}}}\left|S_{\widehat{\epsilon}}\left(i+\lceil\sqrt{n}As_{\ell}\rceil\right)-S_{z}\left(i+\lceil\sqrt{n}As_{\ell}\rceil\right)\right|$
$+\frac{\|\psi\|_{L^{\infty}}}{\sqrt{ns_{\ell}}}\left|S_{\widehat{\epsilon}}\left(i-\lceil\sqrt{n}As_{\ell}\rceil-1\right)-S_{z}\left(i-\lceil\sqrt{n}As_{\ell}\rceil-1\right)\right|$
[…]
For the third term on the right-hand side of (149), if $\psi$ is absolutely continuous, then its derivative $\psi^{\prime}$ exists almost everywhere, and
[…]
$=$ $\frac{1}{n^{1/4}s_{\ell}^{1/2}}\overset{\lceil\sqrt{n}As_{\ell}\rceil}{\underset{j=-\lceil\sqrt{n}As_{\ell}\rceil}{\sum}}\left|\int_{jn^{-1/2}s_{\ell}^{-1}}^{(j+1)n^{-1/2}s_{\ell}^{-1}}\psi^{\prime}(\tau)d\tau\right|$
(150) $\leq$ $\frac{1}{n^{1/4}s_{\ell}^{1/2}}\|\psi^{\prime}\|_{L^{1}}.$
By substituting (150) into (149),
$\sqrt{s_{\ell}}\left|\widetilde{W}_{\widehat{\epsilon}}(t_{i},s_{\ell})-\widetilde{W}_{z}(t_{i},s_{\ell})\right|\leq$ $\frac{1}{\sqrt{n}}\left(\|\psi^{\prime}\|_{L^{1}}+2\|\psi\|_{L^{\infty}}\right)$
(151) $\left(\underset{-\lceil\sqrt{n}As_{\ell}\rceil-1\leq j\leq\lceil\sqrt{n}As_{\ell}\rceil}{\max}\left|S_{\widehat{\epsilon}}\left(i+j\right)-S_{z}\left(i+j\right)\right|\right).$
Since $D^{(1)}\subset D^{(2)}\subset\cdots\subset D^{(d)}$,
$\underset{1\leq i\leq n}{\max}\ \underset{1\leq\ell\leq d}{\max}\ \underset{-\lceil\sqrt{n}As_{\ell}\rceil-1\leq j\leq\lceil\sqrt{n}As_{\ell}\rceil}{\max}\left|S_{\widehat{\epsilon}}\left(i+j\right)-S_{z}\left(i+j\right)\right|$
$=$ $\underset{-\lceil\sqrt{n}As_{d}\rceil\leq h\leq\lceil\sqrt{n}As_{d}\rceil+n}{\max}\left|S_{\widehat{\epsilon}}\left(h\right)-S_{z}\left(h\right)\right|$
(152) $=$ $\underset{h\in I_{n}}{\max}\left|S_{\widehat{\epsilon}}\left(h\right)-S_{z}\left(h\right)\right|,$
where the last equality follows from that $S_{\widehat{\epsilon}}(-\lceil\sqrt{n}As_{d}\rceil)=S_{z}(-\lceil\sqrt{n}As_{d}\rceil)=0$. By (151) and (152), for any $\varepsilon>0$,
$\mathbb{P}\left(\underset{1\leq\ell\leq d}{\max}\ \underset{1\leq i\leq n}{\max}\sqrt{s_{\ell}}\left|\widetilde{W}_{\widehat{\epsilon}}(t_{i},s_{\ell})-\widetilde{W}_{z}(t_{i},s_{\ell})\right|>\varepsilon\right)$
$\leq$ $\mathbb{P}\left(\frac{1}{\sqrt{n}}\left(\|\psi^{\prime}\|_{L^{1}}+2\|\psi\|_{L^{\infty}}\right)\underset{h\in I_{n}}{\max}\left|S_{\widehat{\epsilon}}\left(h\right)-S_{z}\left(h\right)\right|>\varepsilon\right)$
(153) $=$ $\mathbb{P}\left(\underset{h\in I_{n}}{\max}\left|S_{\widehat{\epsilon}}\left(h\right)-S_{z}\left(h\right)\right|>\frac{\varepsilon\sqrt{n}}{\|\psi^{\prime}\|_{L^{1}}+2\|\psi\|_{L^{\infty}}}\right).$
Under Assumptions 3 and 4, Theorem 3.1 in [33] shows that
(154) $\mathbb{E}\left[\underset{h\in I_{n}}{\max}\left|\left(S_{\widehat{\epsilon}}\left(h\right)-S_{z}\left(h\right)\right)\right|^{2}\right]=\mathcal{O}\left(|I_{n}|^{1-G(p,\chi)}\left(\log|I_{n}|\right)^{4}\right),$
where the cardinality of $I_{n}$, denoted by $|I_{n}|$, is equal to $2\lceil\sqrt{n}As_{d}\rceil+n$. By applying the Markov inequality to (153),
$\mathbb{P}\left(\underset{1\leq\ell\leq d}{\max}\ \underset{1\leq i\leq n}{\max}\sqrt{s_{\ell}}\left|\widetilde{W}_{\widehat{\epsilon}}(t_{i},s_{\ell})-\widetilde{W}_{z}(t_{i},s_{\ell})\right|>\varepsilon\right)$
$=$ $\left(\|\psi^{\prime}\|_{L^{1}}+2\|\psi\|_{L^{\infty}}\right)^{2}\varepsilon^{-2}n^{-1}\mathcal{O}\left(|I_{n}|^{1-G(p,\chi)}\left(\log|I_{n}|\right)^{4}\right)$
(155) $=$ $\varepsilon^{-2}\mathcal{O}\left(n^{-G(p,\chi)}\left(\log n\right)^{4}\right).$
By choosing
$\varepsilon=\varepsilon_{n}:=n^{-G(p,\chi)/3}(\log n)^{4/3+\delta},$
where $\delta$ is an arbitrary positive number, (155) becomes
(156) $\mathbb{P}\left(\underset{\begin{subarray}{c}1\leq i\leq n\\ 1\leq\ell\leq d\end{subarray}}{\max}\sqrt{s_{\ell}}|\widetilde{W}_{\widehat{\epsilon}}(t_{i},s_{\ell})-\widetilde{W}_{z}(t_{i},s_{\ell})|>\varepsilon_{n}\hskip-4.26773pt\right)=o\left(\varepsilon_{n}\right)$
when $n\rightarrow\infty$. Denote
$\widetilde{\mathbf{W}}_{\square}=\{\sqrt{s_{\ell}}\widetilde{W}_{\square}(t_{i},s_{\ell})\mid 1\leq i\leq n,\ 1\leq\ell\leq d\},$
where $\square\in\{\epsilon,\widehat{\epsilon},z\}$. For any $B\in\mathcal{B}(\mathbb{C}^{n\times d})$, by (146),
(157) $\nu_{\epsilon}^{(n)}\left(B\right)=$ $\mathbb{P}\left(\widetilde{\mathbf{W}}_{\epsilon}\in B\right)=\mathbb{P}\left(\widetilde{\mathbf{W}}_{\widehat{\epsilon}}\in B\right)$
$=$ $\mathbb{P}\left(\widetilde{\mathbf{W}}_{\widehat{\epsilon}}\in B,|\widetilde{\mathbf{W}}_{\widehat{\epsilon}}-\widetilde{\mathbf{W}}_{z}|_{\infty}\leq\varepsilon_{n}\right)+\mathbb{P}\left(\widetilde{\mathbf{W}}_{\widehat{\epsilon}}\in B,|\widetilde{\mathbf{W}}_{\widehat{\epsilon}}-\widetilde{\mathbf{W}}_{z}|_{\infty}>\varepsilon_{n}\right)$
$\leq$ $\mathbb{P}\left(\widetilde{\mathbf{W}}_{z}\in B^{\varepsilon_{n}}\right)+\mathbb{P}\left(|\widetilde{\mathbf{W}}_{\widehat{\epsilon}}-\widetilde{\mathbf{W}}_{z}|_{\infty}>\varepsilon_{n}\right).$
By (156) and (157),
$\nu_{\epsilon}^{(n)}\left(B\right)<$ $\nu_{z}^{(n)}\left(B^{\varepsilon_{n}}\right)+o\left(\varepsilon_{n}\right)$
[…]
Remark 4.
By (151), the Euclidean norm in $\mathbb{R}^{d}$ of $\widetilde{\mathbf{W}}_{\widehat{\epsilon}}(t_{i})-\widetilde{\mathbf{W}}_{z}(t_{i})$ is bounded as follows.
$\left|\widetilde{\mathbf{W}}_{\widehat{\epsilon}}(t_{i})-\widetilde{\mathbf{W}}_{z}(t_{i})\right|^{2}=\overset{d}{\underset{\ell=1}{\sum}}\left|\widetilde{W}_{\widehat{\epsilon}}(t_{i},s_{\ell})-\widetilde{W}_{z}(t_{i},s_{\ell})\right|^{2}$
$\leq$ $\left(\|\psi^{\prime}\|_{L^{1}}+2\|\psi\|_{L^{\infty}}\right)^{2}n^{-1}\overset{d}{\underset{\ell=1}{\sum}}s_{\ell}^{-1}\hskip-2.84544pt\underset{-\lceil\sqrt{n}As_{\ell}\rceil-1\leq j\leq\lceil\sqrt{n}As_{\ell}\rceil}{\max}\left|S_{\widehat{\epsilon}}\left(i+j\right)-S_{z}\left(i+j\right)\right|^{2},$
which implies that
$\left(\|\psi^{\prime}\|_{L^{1}}+2\|\psi\|_{L^{\infty}}\right)^{-2}n\underset{1\leq i\leq n}{\max}\ \left|\widetilde{\mathbf{W}}_{\widehat{\epsilon}}(t_{i})-\widetilde{\mathbf{W}}_{z}(t_{i})\right|^{2}$
$\leq$ $\overset{d}{\underset{\ell=1}{\sum}}s_{\ell}^{-1}\ \underset{-\lceil\sqrt{n}As_{\ell}\rceil\leq h\leq\lceil\sqrt{n}As_{\ell}\rceil+n}{\max}\left|S_{\widehat{\epsilon}}\left(h\right)-S_{z}\left(h\right)\right|^{2}.$
Hence,
$\left(\|\psi^{\prime}\|_{L^{1}}+2\|\psi\|_{L^{\infty}}\right)^{-2}n\mathbb{E}\left[\underset{1\leq i\leq n}{\max}\ \left|\widetilde{\mathbf{W}}_{\widehat{\epsilon}}(t_{i})-\widetilde{\mathbf{W}}_{z}(t_{i})\right|^{2}\right]$
$\leq$ $\overset{d}{\underset{\ell=1}{\sum}}s_{\ell}^{-1}\mathbb{E}\left[\underset{-\lceil\sqrt{n}As_{\ell}\rceil\leq h\leq\lceil\sqrt{n}As_{\ell}\rceil+n}{\max}\left|S_{\widehat{\epsilon}}\left(h\right)-S_{z}\left(h\right)\right|^{2}\right].$
By applying (154) to the inequality above,
$n\mathbb{E}\left[\underset{1\leq i\leq n}{\max}\left|\widetilde{\mathbf{W}}_{\widehat{\epsilon}}(t_{i})-\widetilde{\mathbf{W}}_{z}(t_{i})\right|^{2}\right]$
$=$ $\overset{d}{\underset{\ell=1}{\sum}}\frac{1}{s_{\ell}}\ \mathcal{O}\left(\left(2\lceil\sqrt{n}As_{\ell}\rceil+n\right)^{1-G(p,\chi)}\left(\log\left(2\lceil\sqrt{n}As_{\ell}\rceil+n\right)\right)^{4}\right),$
which implies that
$\mathbb{E}\left[\underset{1\leq i\leq n}{\max}\left|\widetilde{\mathbf{W}}_{\widehat{\epsilon}}(t_{i})-\widetilde{\mathbf{W}}_{z}(t_{i})\right|^{2}\right]=d\mathcal{O}\left(n^{-G(p,\chi)}\left(\log\left(n\right)\right)^{4}\right)$
if the scale values $\{s_{\ell}\}_{1\leq\ell\leq d}$ are bounded.